# Linear constellations in primes with arithmetic restrictions

### v2

@@ 5.1. Ideal norms and equidistributed nilsequences
We will need cancellation when equidistributed nilsequences of mean zero are summed over the norms of ideals of a number field $K$ (i.e. nonzero ideals of the ring of integers $\mathscr{O}_{K}$), twisted by Dirichlet characters of $K$. A result of Kane [23, Lemma 21] deals with the case of exponential phases $e(\alpha n)$. Here, we generalise Kane’s result to the non-abelian setting.
@@
as then $a_{i}\in\mathfrak{N}_{K/\mathbb{Q}}(\mathfrak{c})\mathscr{O}_{L}\cap\mathbb{Q}=\mathfrak{N}_{K/\mathbb{Q}}(\mathfrak{c})\mathbb{Z}$. To show (5.2), we first consider a prime ideal $\mathfrak{P}$ of $L$. With $p=\mathfrak{P}\cap\mathbb{Z}$ and the factorisation $p\mathscr{O}_{L}=(\mathfrak{P}_{1}\cdots\mathfrak{P}_{r})^{e}$ with $\mathfrak{P}_{1}=\mathfrak{P}$ and $[\mathscr{O}_{L}/\mathfrak{P}_{i}:\mathbb{F}_{p}]=f$, we get
@@
Recall that a Dirichlet character $\xi$ of $K$ is a finite order Hecke character. We consider it as a character of the ray class group $\xi:I(\mathfrak{m})/P_{\mathfrak{m}}\to S^{1}$, where $\mathfrak{m}=\mathfrak{m}_{\infty}\mathfrak{m}_{0}$ is a cycle, $I(\mathfrak{m})$ is the group of fractional ideals relatively prime to $\mathfrak{m}_{0}$ and $P_{\mathfrak{m}}$ is the subgroup of principal fractional ideals $(\alpha)\in I(\mathfrak{m})$ with $\alpha\equiv 1\operatorname{mod*}\mathfrak{m}$. See [26, Chapter VI] for more on this notation. As usual, for arbitrary nonzero ideals $\mathfrak{a}$ of $\mathscr{O}_{K}$, we set $\xi(\mathfrak{a}):=\xi(\mathfrak{a}P_{\mathfrak{m}})$ if $\mathfrak{a}\in I(\mathfrak{m})$ and $\xi(\mathfrak{a}):=0$ otherwise. Here is our version of [23, Lemma 21].
@@
We start with a few standard reductions as in the proof of [15, Proposition 2.1]. As $\lVert F\rVert_{\infty}\leq\lVert F\rVert_{\operatorname{Lip}}$, the desired bound holds trivially whenever $\delta\gg_{m,d,K,\mathfrak{m}}1$. Moreover, if $\delta\leq 1/N$, then the total $\delta$-equidistribution of $(g(n)\Gamma)_{n\in[N]}$ implies that $|\xi(\mathfrak{a})F(g(\mathfrak{N}\mathfrak{a})\Gamma)|\leq\delta\lVert F\rVert_{\operatorname{Lip}}$ for every ideal $\mathfrak{a}$ with $\mathfrak{N}\mathfrak{a}\leq N$, again making the conclusion trivial. Hence, we may assume that $\delta>N^{-1/2}$. Indeed, for $\delta>1/N$ we may then apply the result for $\delta^{1/2}$, thus halving the value of $c(m,d,D)$. Therefore, we assume from now on that
@@
Identifying each $K_{v}$ with $\mathbb{R}$ or $\mathbb{C}$ provides us with notions of euclidean norm and distance, as well as volume on $K_{\infty}$ and thus on $K_{\infty}^{\times}$. Let $V$ be a free abelian subgroup of $U_{\mathfrak{m}}$ that generates $U_{\mathfrak{m}}$ modulo roots of unity. Then $V$ is finitely generated by Dirichlet’s unit theorem. A classical construction (see e.g. [26, VI,§3, Lemma 1]) yields a fundamental domain $\mathscr{F}\subset J_{K}(\infty,\mathfrak{m})$ for the action of $U_{\mathfrak{m}}$ on $J(\infty,\mathfrak{m})$ with the following properties: $t\mathscr{F}=\mathscr{F}$ for all $t>0$, and the set $\mathscr{F}(1)\subseteq\mathscr{F}$ has a $(D-1)$-Lipschitz-parameterisable boundary, where
@@
It follows from these properties that $\mathscr{F}(B)=B^{1/D}\mathscr{F}(1)$, in particular $\mathscr{F}(B)$ has diameter $\asymp_{K,\mathfrak{m}}B^{1/D}$ and a $(D-1)$-Lipschitz-parameterisable boundary with Lipschitz constant $\asymp_{K,\mathfrak{m}}B^{1/D}$.
@@
Recall that $q\ll\delta^{-c/2}$. Hence we may choose $C$ and $N$ sufficiently large in order to ensure that the set $\{n\in[M]\ :\ n\equiv v\bmod q\}$ is a subprogression of $[\tilde{N}]$ of length $\geq\delta^{c}\tilde{N}$. Therefore, total $\delta^{c}$-equidistribution yields the estimate
@@
As at the start of the proof of Proposition 5.2, we may assume that $11/N<\delta<$ 1, as otherwise the bound is either trivial or follows immediately from the total $\delta$-equidistribution of $(g(n)\Gamma)_{n\in[N]}$.
@@ 5.2. Ideal von Mangoldt function and Vaughan’s identity
[…]
@@
The divisor moment bound $\mathbb{E}_{n\leq N}\tau(n)^{k}\ll(\log N)^{2^{k}-1}$ is well known, for references see [12, Lemma C.1]. Splitting dyadically, we see that
@@
Using Lemma 5.6, we deduce the following version of [12, Proposition 4.2] for $\Lambda_{K}$ (twisted by a bounded multiplicative function) in place of $\mu$.
@@
Splitting the interval $[1,N^{1/3}]$ dyadically and applying the pigeonhole principle, we find an integer $1\leq R\leq N^{1/3}$, such that
@@
[…]
@@
$=|\mathbb{E}_{B/2<\mathfrak{N}\mathfrak{b},\mathfrak{N}\mathfrak{b}^{\prime}\leq W}\mathbb{E}_{R/2<r,r^{\prime}\leq R}f(r\mathfrak{N}\mathfrak{b})f(r\mathfrak{N}\mathfrak{b}^{\prime})f(r^{\prime}\mathfrak{N}\mathfrak{b})f(r^{\prime}\mathfrak{N}\mathfrak{b}^{\prime})|.$
@@
We follow the proof of [14, Proposition 2.1], replacing the application of [14, Proposition 3.1] by our Propositions 5.7 and 5.2.
@@
Hence, there is a progression $\tilde{P}_{r}=\{a+q,a+2q,\ldots,a+N_{r}q\}\subseteq[M_{r}]$ of length $N_{r}\geq\delta^{c_{1}}M_{r}$, such that the sequence $(g_{r}(a+nq)\Gamma)_{n\in[N_{r}]}$ is not $\delta^{c_{1}}$-equidistributed. Define the polynomial sequence $\tilde{g}_{r}(n):=g_{r}(a+nq)$. By the quantitative Leibmann theorem of Green and Tao [15, Theorem 2.9], there exists a nontrivial horizontal character $\psi_{r}$ with $|\psi_{r}|\ll\delta^{-O_{m,d}(c_{1})}$, such that
@@
Hence, we may assume that $R\delta^{O_{D}(c_{1})}\gg 1$ and, replacing $R$ by $2\lfloor R/2\rfloor$, that $R$ is even. Now the arguments in [14, pp.550–551] apply verbatim with $K=R/2$ in their notation, leading to the same contradiction.
@@
We can therefore find an arithmetic progression $\tilde{P}_{r,r^{\prime}}=\{a+q,\ldots,a+N_{r,r^{\prime}}q\}\subseteq[B]$ with length $N_{r,r^{\prime}}\geq\delta^{c_{1}}B$ so that $(g_{r,r^{\prime}}(a+nq)(\Gamma\times\Gamma))_{n\in[N_{r,r^{\prime}}]}$ fails to be totally $\delta^{c_{1}}$-equidistributed. With $\tilde{g}_{r,r^{\prime}}(n):=g_{r,r^{\prime}}(a+nq)$, the quantitative Leibmann theorem [15, Theorem 2.9] yields the existence of a nontrivial horizontal character $\psi_{r,r^{\prime}}$ with $|\psi_{r,r^{\prime}}|\ll\delta^{-O_{m,d}(c_{1})}$, such that
@@
If $C$ is large enough, this holds for $\gg R^{2}\delta^{O_{D}(c)}\geq R^{2}\delta^{O_{D}(c_{1})}$ values of $(r,r^{\prime})$ with $R/2<r,r^{\prime}\leq R$. As $R\gg N^{1/3}$, we have $R^{2}\delta^{O_{D}(c_{1})}\gg R$, and hence we may assume that $R$ is even. From here on, the proof in [14, pp.553] applies verbatim with $K=R/2$. ∎
@@
Let $10\leq l\leq\lfloor-(1/2)\log_{2}\delta\rfloor-$ 1, let $R\subseteq[N_{l}]\subseteq[N]$ be an arithmetic progression of length at least $\delta^{1/2}N_{l}$. As $\delta^{1/2}N_{l}\geq\delta^{1/2}x_{l}\geq\delta N$, the total $\delta$-equidistribution of $(g(n)\Gamma)_{n\in[N]}$ yields
@@
In this section, we prove Theorem 2.3. As in [14], we deduce the general result from the following version in the case of equidistributed nilsequences and test functions with mean zero. This result is analogous to [14, Proposition 2.1], with the Möbius function replaced by our $W$-tricked Chebotarev-von Mangoldt function $\Lambda_{K,C,b,W}(n)$.
@@
Let $N\in\mathbb{N}$ be sufficiently large in terms of $m,d$. Let $\delta\in(0,1)$ and $Q\geq 2$. Let $G/\Gamma$ be an $m$-dimensional nilmanifold with a filtration $G_{\bullet}$ of degree $d$ and a $Q$-rational Mal’cev basis $\mathscr{X}$ adapted to $G_{\bullet}$. Let $g\in\operatorname{poly}(\mathbb{Z},G_{\bullet})$, and suppose that $(g(n)\Gamma)_{n\in[N]}$ is totally $\delta$-equidistributed. Let $F:G/\Gamma\to[-1,1]$ with $\int_{G/\Gamma}F=0$ and $P\subseteq[N]$ an arithmetic progression of size $|P|\geq N/Q$. Let $K$ be a Galois number field of degree $[K:\mathbb{Q}]=D$ with $\Phi_{K^{\operatorname{ab}}}\mid W$, and $C\subseteq\operatorname{Gal}(K/\mathbb{Q})$ a conjugacy class. Then
$\left|\mathbb{E}_{n\in[N]}\Lambda_{K,C,b,W}(n)\mathbf{1}_{P}(n)F(g(n)\Gamma)\right|\ll_{K,m,d}\delta^{c(m,d,D)}Q\lVert F\rVert_{\operatorname{Lip}}(\log N)^{4}.$
@@ 6.1. Passing to ideals
For our Galois number field $K$ and conjugacy class $C\subseteq\operatorname{Gal}(K/\mathbb{Q})$, fix in addition an element $c\in C$. We consider the fixed field $L=K^{c}$. Then $K/L$ is cyclic with Galois group $\operatorname{Gal}(L/K)=\langle c\rangle$. We consider every character $\xi:\langle c\rangle\to S^{1}$ as a Hecke character of $L$ via $\xi(\mathfrak{p})=\xi([K/L,\mathfrak{p}])$ for all prime ideals $\mathfrak{p}$ of $L$ unramified in $K/L$. The following lemma and its proof are inspired by [23, Proposition 6].
@@
Let $p$ be a prime unramified in $K$ with $[K/\mathbb{Q},p]=C$. The Galois group $\operatorname{Gal}(K/\mathbb{Q})$ acts transitively on the prime ideals of $\mathscr{O}_{K}$ lying above $p$, the stabilisers being the decomposition groups, which have size $|c|$. Hence, there are exactly $[K:\mathbb{Q}]/|c|$ prime ideals $\mathfrak{q}$ of $\mathscr{O}_{K}$ above $p$, of which exactly $[K:\mathbb{Q}]/(|c|\cdot|C|)$ satisfy $[K/\mathbb{Q},\mathfrak{q}]=c$.
@@
$\ll_{m,d,L}\delta^{c(m,d)c(m,d,D_{L})}QW\lVert F\rVert_{\operatorname{Lip}}W(N+1)\log(W(N+1))^{2}\ll\delta^{c(m,d)c(m,d,D_{L})}Q\lVert F\rVert_{\operatorname{Lip}}N(\log N)^{4},$

### v3

@@ 5.1. Ideal norms and equidistributed nilsequences
We will need cancellation when equidistributed nilsequences of mean zero are summed over the norms of ideals of a number field $K$ (i.e. nonzero ideals of the ring of integers $\mathscr{O}_{K}$), twisted by Dirichlet characters of $K$. A result of Kane [24, Lemma 21] deals with the case of exponential phases $e(\alpha n)$. Here, we generalise Kane’s result to the non-abelian setting.
@@
as then $a_{i}\in\mathfrak{N}_{K/\mathbb{Q}}(\mathfrak{c})\mathscr{O}_{L}\cap\mathbb{Q}=\mathfrak{N}_{K/\mathbb{Q}}(\mathfrak{c})\mathbb{Z}$. To show (5.2), we first consider a prime ideal $\mathfrak{P}$ of $L$. With $p=\mathfrak{P}\cap\mathbb{Z}$ and the factorisation $p\mathscr{O}_{L}=(\mathfrak{P}_{1}\cdots\mathfrak{P}_{r})^{e}$ with $\mathfrak{P}_{1}=\mathfrak{P}$ and $\mathscr{O}_{L}/\mathfrak{P}_{i}:\mathbb{F}_{p}]=f$, we get
@@
Recall that a Dirichlet character $\xi$ of $K$ is a finite order Hecke character. We consider it as a character of the ray class group $\xi:I(\mathfrak{m})/P_{\mathfrak{m}}\to S^{1}$, where $\mathfrak{m}=\mathfrak{m}_{\infty}\mathfrak{m}_{0}$ is a cycle, $I(\mathfrak{m})$ is the group of fractional ideals relatively prime to $\mathfrak{m}_{0}$ and $P_{\mathfrak{m}}$ is the subgroup of principal fractional ideals $(\alpha)\in I(\mathfrak{m})$ with $\alpha\equiv 1\operatorname{mod*}\mathfrak{m}$. See [28, Chapter VI] for more on this notation. As usual, for arbitrary nonzero ideals $\mathfrak{a}$ of $\mathscr{O}_{K}$, we set $\xi(\mathfrak{a}):=\xi(\mathfrak{a}P_{\mathfrak{m}})$ if $\mathfrak{a}\in I(\mathfrak{m})$ and $\xi(\mathfrak{a}):=0$ otherwise. Here is our version of [24, Lemma 21].
@@
We start with a few standard reductions as in the proof of [16, Proposition 2.1]. As $\lVert F\rVert_{\infty}\leq\lVert F\rVert_{\operatorname{Lip}}$, the desired bound holds trivially whenever $\delta\gg_{m,d,K,\mathfrak{m}}1$. Moreover, if $\delta\leq 1/N$, then the total $\delta$-equidistribution of $(g(n)\Gamma)_{n\in[N]}$ implies that $|\xi(\mathfrak{a})F(g(\mathfrak{N}\mathfrak{a})\Gamma)|\leq\delta\lVert F\rVert_{\operatorname{Lip}}$ for every ideal $\mathfrak{a}$ with $\mathfrak{N}\mathfrak{a}\leq N$, again making the conclusion trivial. Hence, we may assume that $\delta>N^{-1/2}$. Indeed, for $\delta>1/N$ we may then apply the result for $\delta^{1/2}$, thus halving the value of $c(m,d,D)$. Therefore, we assume from now on that
@@
Identifying each $K_{v}$ with $\mathbb{R}$ or $\mathbb{C}$ provides us with notions of euclidean norm and distance, as well as volume on $K_{\infty}$ and thus on $K_{\infty}^{\times}$. Let $V$ be a free abelian subgroup of $U_{\mathfrak{m}}$ that generates $U_{\mathfrak{m}}$ modulo roots of unity. Then $V$ is finitely generated by Dirichlet’s unit theorem. A classical construction (see e.g. [28, VI,§3, Lemma 1]) yields a fundamental domain $\mathscr{F}\subset J_{K}(\infty,\mathfrak{m})$ for the action of $U_{\mathfrak{m}}$ on $J(\infty,\mathfrak{m})$ with the following properties: $t\mathscr{F}=\mathscr{F}$ for all $t>0$, and the set $\mathscr{F}(1)\subseteq\mathscr{F}$ has a (D1)(D-1)-Lipschitz-parameterisable boundary, where
@@
It follows from these properties that $\mathscr{F}(B)=B^{1/D}\mathscr{F}(1)$, in particular $\mathscr{F}(B)$ has diameter $\asymp_{K,\mathfrak{m}}B^{1/D}$ and a (D1)(D-1)-Lipschitz-parameterisable boundary with Lipschitz constant $\asymp_{K,\mathfrak{m}}B^{1/D}$.
@@
Recall that $q\ll\delta^{-c/2}$. Hence we may choose $C$ and $N$ sufficiently large in order to ensure that the set $\{n\in[M]\ :\ n\equiv v\bmod q\}$ is a subprogression of $\tilde{N}$] of length $\geq\delta^{c}\tilde{N}$. Therefore, total $\delta^{c}$-equidistribution yields the estimate
@@
As at the start of the proof of Proposition 5.2, we may assume that $1/N<\delta<1$, as otherwise the bound is either trivial or follows immediately from the total $\delta$-equidistribution of $(g(n)\Gamma)_{n\in[N]}$.
@@ 5.2. Ideal von Mangoldt function and Vaughan’s identity
$a_{K}(n)$ $:=\#\{\mathfrak{a}\text{ ideal of }\mathscr{O}_{K}\ :\ \mathfrak{N}\mathfrak{a}=n\},$
@@
The divisor moment bound $\mathbb{E}_{n\leq N}\tau(n)^{k}\ll(\log N)^{2^{k}-1}$ is well known, for references see [13, Lemma C.1]. Splitting dyadically, we see that
@@
Using Lemma 5.6, we deduce the following version of [13, Proposition 4.2] for $\Lambda_{K}$ (twisted by a bounded multiplicative function) in place of $\mu$.
@@
Splitting the interval $1,N^{1/3}]$ dyadically and applying the pigeonhole principle, we find an integer $1\leq R\leq N^{1/3}$, such that
@@
with some coefficients $\mathbf{b}(\mathfrak{b},\mathfrak{b}^{\prime})\in\mathbb{C}$ that satisfy $|\mathbf{b}(\mathfrak{w},\mathfrak{w}^{\prime})|\leq 1$. Applying [13, Lemma A.10] with $x=(\mathfrak{b},\mathfrak{b}^{\prime})$, $y=r$, $X=\{(\mathfrak{b},\mathfrak{b}^{\prime})\ :\ \mathfrak{b},\mathfrak{b}^{\prime}\in\mathscr{I}_{K},\ B/2<\mathfrak{N}\mathfrak{b},\mathfrak{N}\mathfrak{b}^{\prime}\leq B\}$, $Y=\mathbb{Z}\cap(R/2,R]$ and $h(x,y)=f(r\mathfrak{N}\mathfrak{b})f(r\mathfrak{N}\mathfrak{b}^{\prime})$, we obtain
@@
$\hskip-56.9055pt=|\mathbb{E}_{B/2<\mathfrak{N}\mathfrak{b},\mathfrak{N}\mathfrak{b}^{\prime}\leq W}\mathbb{E}_{R/2<r,r^{\prime}\leq R}f(r\mathfrak{N}\mathfrak{b})f(r\mathfrak{N}\mathfrak{b}^{\prime})f(r^{\prime}\mathfrak{N}\mathfrak{b})f(r^{\prime}\mathfrak{N}\mathfrak{b}^{\prime})|.$
@@
We follow the proof of [15, Proposition 2.1], replacing the application of [15, Proposition 3.1] by our Propositions 5.7 and 5.2.
@@
Hence, there is a progression $\tilde{P}_{r}=\{a+q,a+2q,\ldots,a+N_{r}q\}\subseteq[M_{r}]$ of length $N_{r}\geq\delta^{c_{1}}M_{r}$, such that the sequence $(g_{r}(a+nq)\Gamma)_{n\in[N_{r}]}$ is not $\delta^{c_{1}}$-equidistributed. Define the polynomial sequence $\tilde{g}_{r}(n):=g_{r}(a+nq)$. By the quantitative Leibmann theorem of Green and Tao [16, Theorem 2.9], there exists a nontrivial horizontal character $\psi_{r}$ with $|\psi_{r}|\ll\delta^{-O_{m,d}(c_{1})}$, such that
@@
Hence, we may assume that $R\delta^{O_{D}(c_{1})}\gg 1$ and, replacing $R$ by $2\lfloor R/2\rfloor$, that $R$ is even. Now the arguments in [15, pp.550–551] apply verbatim with $K=R/2$ in their notation, leading to the same contradiction.
@@
We can therefore find an arithmetic progression $\tilde{P}_{r,r^{\prime}}=\{a+q,\ldots,a+N_{r,r^{\prime}}q\}\subseteq[B]$ with length $N_{r,r^{\prime}}\geq\delta^{c_{1}}B$ so that $(g_{r,r^{\prime}}(a+nq)(\Gamma\times\Gamma))_{n\in[N_{r,r^{\prime}}]}$ fails to be totally $\delta^{c_{1}}$-equidistributed. With $\tilde{g}_{r,r^{\prime}}(n):=g_{r,r^{\prime}}(a+nq)$, the quantitative Leibmann theorem [16, Theorem 2.9] yields the existence of a nontrivial horizontal character $\psi_{r,r^{\prime}}$ with $|\psi_{r,r^{\prime}}|\ll\delta^{-O_{m,d}(c_{1})}$, such that
@@
If $C$ is large enough, this holds for $\gg R^{2}\delta^{O_{D}(c)}\geq R^{2}\delta^{O_{D}(c_{1})}$ values of $(r,r^{\prime})$ with $R/2<r,r^{\prime}\leq R$. As $R\gg N^{1/3}$, we have $R^{2}\delta^{O_{D}(c_{1})}\gg R$, and hence we may assume that $R$ is even. From here on, the proof in [15, pp.553] applies verbatim with $K=R/2$. ∎
@@
Let $0\leq l\leq\lfloor-(1/2)\log_{2}\delta\rfloor-1$, let $R\subseteq[N_{l}]\subseteq[N]$ be an arithmetic progression of length at least $\delta^{1/2}N_{l}$. As $\delta^{1/2}N_{l}\geq\delta^{1/2}x_{l}\geq\delta N$, the total $\delta$-equidistribution of $(g(n)\Gamma)_{n\in[N]}$ yields
@@
In this section, we prove Theorem 2.3. As in [15], we deduce the general result from the following version in the case of equidistributed nilsequences and test functions with mean zero. This result is analogous to [15, Proposition 2.1], with the Möbius function replaced by our $W$-tricked Chebotarev-von Mangoldt function $\Lambda_{K,C,b,W}(n)$.
@@
Let $N\in\mathbb{N}$ be sufficiently large in terms of $m,d$. Let $\delta\in(0,1)$ and $Q\geq 2$. Let $G/\Gamma$ be an $m$-dimensional nilmanifold with a filtration $G_{\bullet}$ of degree $d$ and a $Q$-rational Mal’cev basis $\mathscr{X}$ adapted to $G_{\bullet}$. Let $g\in\operatorname{poly}(\mathbb{Z},G_{\bullet})$, and suppose that $(g(n)\Gamma)_{n\in[N]}$ is totally $\delta$-equidistributed. Let $F:G/\Gamma\to[-1,1]$ with $\int_{G/\Gamma}F=0$ and $P\subseteq[N]$ an arithmetic progression of size $|P|\geq N/Q$. Let $K$ be a Galois number field of degree $K:\mathbb{Q}]=D$ with $\Phi_{K^{\operatorname{ab}}}\mid W$, and $C\subseteq\operatorname{Gal}(K/\mathbb{Q})$ a conjugacy class. Then
$\left|\mathbb{E}_{n\in[N]}\Lambda_{K,C,b,W}(n)\mathbf{1}_{P}(n)F(g(n)\Gamma)\right|\ll_{K,m,d,\mathscr{D}}\delta^{c(m,d,D)}Q\lVert F\rVert_{\operatorname{Lip}}(\log N)^{4}.$
@@ 6.1. Passing to ideals
For our Galois number field $K$ and conjugacy class $C\subseteq\operatorname{Gal}(K/\mathbb{Q})$, fix in addition an element $c\in C$. We consider the fixed field $L=K^{c}$. Then $K/L$ is cyclic with Galois group $\operatorname{Gal}(L/K)=\langle c\rangle$. We consider every character $\xi:\langle c\rangle\to S^{1}$ as a Hecke character of $L$ via $\xi(\mathfrak{p})=\xi([K/L,\mathfrak{p}])$ for all prime ideals $\mathfrak{p}$ of $L$ unramified in $K/L$. The following lemma and its proof are inspired by [24, Proposition 6].
@@
Let $p$ be a prime unramified in $K$ with $K/\mathbb{Q},p]=C$. The Galois group $\operatorname{Gal}(K/\mathbb{Q})$ acts transitively on the prime ideals of $\mathscr{O}_{K}$ lying above $p$, the stabilisers being the decomposition groups, which have size $|c|$. Hence, there are exactly $K:\mathbb{Q}]/|c|$ prime ideals $\mathfrak{q}$ of $\mathscr{O}_{K}$ above $p$, of which exactly $K:\mathbb{Q}]/(|c|\cdot|C|)$ satisfy $K/\mathbb{Q},\mathfrak{q}]=c$.
@@
$\ll_{m,d,L}\delta^{c(m,d)c(m,d,D_{L})}QW\lVert F\rVert_{\operatorname{Lip}}W(N+1)\log(W(N+1))^{2}\ll_{\mathscr{D}}\delta^{c(m,d)c(m,d,D_{L})}Q\lVert F\rVert_{\operatorname{Lip}}N(\log N)^{4},$